Status Report

SOHO Daily Meeting Minutes for Tuesday, 11 June 2002

By SpaceRef Editor
June 11, 2002
Filed under , ,

* RHESSI target for today is AR 9987; target for tomorrow
is TBD.

* D46 data from DOY 162 02:13 to 03:37 UT did not arrive this
morning. It has been requested.

* We may lose Madrid contacts on 18 & 19 June due to a strike
in Spain.

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, New tracking star, mag = 6.35

Planned Activities: VIRGO, SWAN, CELIAS, RSL

Upcoming Operations: None

Ground Anomalies:
161/1300 AOS from D27 seven minutes late due to LMC failure.
Handover activities delayed seven minutes; no TLM
or CMD loss due to coverage from D66.
161/1719 Received degraded TLM form 1719 to 1725 due to RFI.
Any lost data is unrecoverable.

06/10 DOY 161 – 06/11 DOY 162

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. AR/flare diagnostics on AR9991 and JOP146.
Tomorrow: ICAL, JOP146, sunspot study on AR9991,

UVCS: Nominal. Coronal hole observations and synoptic.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. 195 CME watch and synoptic, ICAL tomorrow.

MDI: Nominal. Full disk magnetograms and dopplergrams.

TRACE: Nominal. JOP146 targetting AR 9978 and AR9773/9775.
Tomorrow: JOP146 with the same target area.

SpaceRef staff editor.